# Investigating dusty Red Supergiant outflows in Westerlund 1 with 3D Hydrodynamic simulations

### v1

@@ Abstract
Context. Recent JWST observations towards Westerlund 1 revealed extensive nebular emission associated with the cluster. Given the age of the region and proximity of that material to massive stars it cannot be primordial star forming gas and the origin is uncertain.
Aims. We aim to determine whether the nebular emission in Westerlund 1 could be due to ablation flows from Red Supergiant (RSG) stars embedded in the cluster wind driven by the Wolf-Rayet stars in the cluster core. We also aim to explore the efficiency of mass-loading for the RSG wind in this scenario.
Methods. We use 3D hydrodynamic simulations with the pion code to study the interaction between the cluster and RSG winds. We compare with the JWST observations by generating synthetic dust-emission maps.
Results. We find that the ablation flow morphology is consistent with the observations towards Westerlund 1, with clumps and instabilities. Synthetic observations at 11 m predict fluxes in the ablation flow of MJy ster-1 which is consistent with the unsaturated components of the JWST F1130W observations in the vicinity of the red supergiants in the region. This good agreement is achieved without any consideration of polycyclic aromatic hydrocarbons (PAHs), which have a known 11.3 m feature that appears in the F1130W band. This suggests that the ablation flow is PAH depleted.
@@
Young massive stellar clusters (YMCs) provide ideal environments to study the evolution of massive stars, planet and star formation processes and stellar feedback to the Galaxy (Portegies Zwart et al., 2010). In recent times, they have also been proposed as sources of cosmic rays due to detections of gamma-rays (e.g. H.E.S.S. Collaboration et al., 2022). Westerlund 1 is thought to be the most massive YMC in the Galaxy, with mass estimates on the order of to (Clark et al., 2005; Andersen et al., 2017), although the inferred mass is dependent on the derived luminosity (hence age) of the massive stars in the cluster, which is quite uncertain because of the large extinction to the cluster (Beasor et al., 2021). Westerlund 1 hosts a large and diverse population of massive stars at various evolutionary stages, with 24 Wolf-Rayet (WR) stars (Crowther et al., 2006) as well as OB supergiants, yellow supergiants (YSG), red supergiants (RSG), Luminous Blue Variables and an sgB[e] star (Clark et al., 2020). The WR stars drive a cluster wind of several 1000 km s-1 (Härer et al., 2023) which sculpts a superbubble cavity around the cluster. The region around Westerlund 1 is a bright gamma-ray emitter to very high energies (H.E.S.S. Collaboration et al., 2022), likely due to the combined action of stellar winds and supernovae in accelerating cosmic rays (e.g. Vieu & Reville, 2023; Härer et al., 2023).
@@
Mass loading of a fast wind by hydrodynamical mixing with cold and slow material was introduced by Hartquist et al. (1986), who applied it to the expansion of WR nebulae and the consequent X-ray and nebular emission that should arise from their model. Mass loading in stellar clusters was considered theoretically by Stevens & Hartwell (2003) in the context of protostars and protoplanetary disks and their effects on diffuse X-ray emission. The effects of stellar winds and supernovae on molecular material remaining from the formation of a YMC were modeled by Rogers & Pittard (2013), who found the cluster wind to be mass loaded by a factor of up to several hundred. Arthur (2012) modelled the X-ray emission from mass loading of protoplanetary disks in Orion. Mass loading of the cluster wind by the dense and slow winds of cool supergiants reduces the thermal pressure driving the cluster wind (Silich et al., 2011), potentially limiting its effectiveness in creating a superbubble and accelerating cosmic rays.
@@
Observations of diffuse thermal X-ray emission in Westerlund 1 show a hard component (Kavanagh et al., 2011; Haubner et al., 2025). While pre-main sequence stars are thought to contribute to this (Clark et al., 2008), the majority is attributed to thermalisation of the winds from WR stars in the cluster core. These winds must mix with cooler material to explain the gas temperature derived from X-ray observations, but the origin of this material is not clear. Dust and ejecta from supernovae are thought to not contribute significantly to this process (Kavanagh et al., 2011; Haubner et al., 2025), since the supernova rate of 1 per kyr would imply the cluster is free of supernova remnant (SNR) material for of the time (Muno et al., 2006), although we note that this estimated supernova rate is very uncertain. RSG winds inject cool material that can be ionized by an external radiation field, such as that found in a YMC (Mackey et al., 2014). This has already been demonstrated for W26, one of the four RSGs in Westerlund 1 (Wright et al., 2014; Mackey et al., 2015). Furthermore, the potential for the enriched RSG wind material to contribute to future star formation within the cluster depends on whether it is heated and ejected rapidly, or can remain confined, cool and dense (Mackey et al., 2015).
@@
Recent JWST images of Westerlund 1 from the Extended Westerlund 1 and 2 Open Clusters Survey (EWOCS) show resolved shells and outflows from the four RSGs and other evolved cool massive stars in the cluster (Guarcello et al., 2024). The emission mechanism for gas and dust in Westerlund 1 is not clear, but its proximity to the core of such a massive cluster implies that the material was recently ejected from the cluster member stars. If the mid-IR emission is produced by dust, then the most likely source is winds and mass ejections from the cool supergiants and hypergiants in the cluster core.
@@
In this paper, we explore in a sense the opposite case to that of Rogers & Pittard (2013), that of a cool RSG wind embedded in a hot, fast collective cluster wind from a YMC. We present a 3D hydrodynamic simulation of our scenario to explore the efficiency of mass loading from evolved stellar winds, and use it to create synthetic dust-emission maps. There is some similarity between our work and previous simulations of the G2 cloud orbiting the Galactic Centre (De Colle et al., 2014; Ballone et al., 2018). They modeled a cool star, emitting a slow wind, moving rapidly through the hot coronal gas around the supermassive black hole Sgr A∗. To our knowledge, no multi-dimensional simulations have been made of the interaction of stellar wind from a RSG embedded in a hot star-cluster wind, although such a situation must be common in YMCs.
@@
Our paper is organised as follows: In Sect. 2, we discuss our simulation setup and values adopted for the RSG and cluster winds. In Sect. 3 we present the results of our simulation, including slices through the XY domain (Sect. 3.1), phase diagrams for the RSG wind and cluster wind material (Sect. 3.2), mass loading of the cluster wind over time (Sect. 3.3), synthetic dust continuum-emission maps for 11 m and 24 m derived from our simulation (Sect. 3.4), and how those maps correspond to observations from JWST (Sect. 3.5). In Sect. 4 we discuss the uncertainties in the winds of RSGs and clusters (Sect. 4.1) and in our modelling (Sect. 4.2), as well as the implications for this work in understanding the role of mass loading (Sect. 4.3) and its connection to Super star-clusters (SSCs) (Sect. 4.4) and the need for follow-up observations (Sect. 4.5). We present our conclusions in Sect. 5, and a resolution study in Appendix A.
@@ 2.1 HD simulations
We perform our 3D Cartesian simulations using the (M)HD code pion (Mackey et al., 2021). We employ the static mesh refinement option in pion to achieve high resolution only where required. We use model 8 in pion as described in Green et al. (2019) for our radiative heating and cooling prescription, which assumes a sufficiently intense EUV radiation field to keep H and He photoionized everywhere on the domain, and which includes three cooling sources. The first source is the maximum of the Solar metallicity collisional ionization equilibrium (CIE) curve described by Wiersma et al. (2009), and the cooling function for forbidden lines described by Henney et al. (2009). This ensures we account for the significant rates of cooling that occur in photoionization equilibrium at K. The second source included is ionized hydrogen (Hummer, 1994) and helium (Rybicki & Lightman, 1979) Bremsstrahlung. The third source is the cooling rate of hydrogen recombination from Hummer (1994), where we assume hydrogen is fully ionized. Heating is calculated by assuming each H recombination (excluding recombinations directly to the ground state, i.e., the Case B approximation) rapidly results in a re-ionization by the intense Lyman-continuum radiation field found in YMCs, with associated photoheating of 5 eV per photoionization. We obtain an equilibrium gas temperature K for this prescription.
@@
We perform our simulation using a 3D Cartesian domain of cm, using 5 static mesh refinement levels. At the coarsest level, , the entire domain is covered, and more refined levels with cover a sub-domain of diameter , i.e., each finer level has a domain smaller than the level preceding it. We include grid cells per level of refinement, such that the maximum grid resolution is cm.
@@
To simulate the stellar wind, we place the RSG star at the origin, and the nested grids are centred on the star. Our aim is to study the ablation of the RSG wind, and thus a constant mass loss source term is applied, with a surface temperature K, a mass-loss rate = 2 -1 (Mackey et al., 2015) and wind terminal velocity (Mauron & Josselin, 2011). The wind injection region is cm, or 10 cells. This is much larger than the stellar radius, and thus we inject the wind at the terminal velocity.
@@
To simulate the cluster wind, we impose a constant x-velocity cm s-1, density g cm-3 and pressure dyn cm-2. While direct observational constraints are challenging, the cluster wind velocities cannot significantly exceed the adiabatic sound speed. This is of order 500-1000 km s-1, using the diffuse X-ray gas temperatures measured by Kavanagh et al. (2011). Our velocity is also consistent with that of Silich et al. (2011), but lower than theoretical predictions (e.g., Stevens & Hartwell, 2003; Härer et al., 2023), which are of order 2-3000 km s-1. The upstream boundary condition was set to the fixed inflow conditions of the cluster wind, and the other external boundaries are set to zero-gradient (outflow).
@@ 3.2 Phase Diagrams
In Fig.2 we show the density-temperature phase diagrams for the RSG wind, cluster wind, and the sum of these components at the end of our simulation. The horizontal part in the lower right of the left panel of Fig. 2 corresponds to material immediately surrounding the star, mainly unshocked stellar wind. The rest of the material is carried away in the cluster wind, where the majority of it remains cool. However, there is a significant tail of RSG gas that is heated and becomes part of the hot phase. In the centre panel there is a “hotspot” at the ambient ISM values, but we see the RSG wind has also cooled some of this material down to the photoionization equilibrium temperature of K. This reflects the dynamical mixing of the two gas phases, which both heat the RSG wind and cool the cluster wind.
In Fig. 3, we show phase diagrams of vs temperature for the stellar wind material at 50 kyr intervals. The wind material is rapidly accelerated as it is heated and then exits the domain, and the effect increases with time. The similarity of the last two panels implies that the simulation is approaching a time-stationary state.
@@ 3.3 Mass loading
To get an idea of the efficiency of the mass loading over time, we compare the time evolution of “hot” (K) and “cool” (K) RSG wind material in Fig. 4. As expected, the cool RSG wind mass increases rapidly at early times. It stays relatively constant over time for small radii, but appears to peak and then decrease for larger radii (Fig. 4a). As the contiguous tail is broken up and ablated, the cool material is removed from the domain more rapidly over time. The mass of hot RSG wind material changes rapidly at first (Fig. 4b), likely due to initial conditions. At later times it tends to increase to a quasi steady state for all radii. As this material is accelerated and exits the domain, we only find a lower limit for the amount of RSG material that is heated. This behaviour is consistent across different heating temperatures at small and large radii (Fig. 4c). The small fluctuations arise from the stochastic nature of the instabilities themselves.
The ratio of hot to cool RSG wind material tends to increase over time, again with stochastic fluctuations. Given that some of the heated wind is escaping from the domain, this represents a lower limit on the mass loading. Taking , this represents a mass-loading efficiency of 33% at late times, albeit a more detailed numerical study with a larger simulation domain is required to measure this number accurately. We study the numerical convergence of these results in Appendix A.

### v2

@@ Abstract
Context. Recent JWST observations towards Westerlund 1 have revealed extensive nebular emission associated with the cluster. Given the age of the region and the proximity of that material to massive stars, it cannot be primordial star-forming gas and the origin is uncertain.
Aims. We aim to determine whether the nebular emission in Westerlund 1 is due to ablation flows from red supergiant (RSG) stars embedded in the cluster wind driven by the Wolf-Rayet stars in the cluster core. We also aim to explore the efficiency of mass loading for the RSG wind in this scenario.
Methods. We used 3D hydrodynamic simulations with the pion code to study the interaction between the cluster and RSG winds. We compared our simulations with the JWST observations by generating synthetic dust-emission maps.
Results. We find that the ablation flow morphology – which shows clumps and instabilities – is consistent with the observations towards Westerlund 1. Synthetic observations at 11 m predict fluxes in the ablation flow of MJy ster-1, which is consistent with the unsaturated components of the JWST F1130W observations in the vicinity of the RSGs in the region. This good agreement is achieved without any consideration of polycyclic aromatic hydrocarbons (PAHs), which have a known 11.3 m feature that appears in the F1130W band. This suggests that the environment is not conducive to PAH formation and/or the ablation flow is PAH-depleted by wind and radiation action.
@@
Young massive stellar clusters (YMCs) provide ideal environments to study the evolution of massive stars, planet and star formation processes, and stellar feedback to the Galaxy (Portegies Zwart et al., 2010). In recent times, they have also been proposed as sources of cosmic rays due to detections of gamma rays (e.g. H.E.S.S. Collaboration et al., 2022). Westerlund 1 is thought to be the most massive YMC in the Galaxy, with mass estimates of the order of to (Clark et al., 2005; Andersen et al., 2017), although the inferred mass is dependent on the derived luminosity (and hence the age) of the massive stars in the cluster, which is quite uncertain because of the large extinction to the cluster (Beasor et al., 2021). Westerlund 1 hosts a large and diverse population of massive stars at various evolutionary stages, with 24 Wolf-Rayet (WR) stars (Crowther et al., 2006) as well as OB supergiants, yellow supergiants, red supergiants (RSGs), luminous blue variables, and an sgB[e] star (Clark et al., 2020). The WR stars drive a cluster wind of several thousand km s-1 (Härer et al., 2023) that sculpts a superbubble cavity around the cluster. The region around Westerlund 1 is a bright gamma-ray emitter to very high energies (H.E.S.S. Collaboration et al., 2022), likely due to the combined action of stellar winds and supernovae in accelerating cosmic rays (e.g. Vieu & Reville, 2023; Härer et al., 2023).
@@
The concept of mass loading of a fast wind by hydrodynamic (HD) mixing with cold and slow material was introduced by Hartquist et al. (1986), who applied it to the expansion of WR nebulae and the consequent X-ray and nebular emission that should arise according to their model. Mass loading in stellar clusters was considered theoretically by Stevens & Hartwell (2003) in the context of protostars and protoplanetary disks and their effects on diffuse X-ray emission. The effects of stellar winds and supernovae on molecular material remaining from the formation of a YMC were modelled by Rogers & Pittard (2013), who found the cluster wind to be mass loaded by a factor of up to several hundred. Arthur (2012) modelled the X-ray emission from the mass loading of protoplanetary disks in Orion. Mass loading of the cluster wind by the dense and slow winds of cool supergiants reduces the thermal pressure driving the cluster wind (Silich et al., 2011), potentially limiting its effectiveness in creating a superbubble and accelerating cosmic rays.
@@
Observations of diffuse thermal X-ray emission in Westerlund 1 show a hard component (Kavanagh et al., 2011; Haubner et al., 2025). While pre-main-sequence stars are thought to contribute to this (Clark et al., 2008), the majority is attributed to thermalisation of the winds from WR stars in the cluster core. These winds must mix with cooler material to explain the gas temperature derived from X-ray observations, but the origin of this material is not clear. Dust and ejecta from supernovae are thought to not contribute significantly to this process (Kavanagh et al., 2011; Haubner et al., 2025) since the supernova rate of 1 per kyr would imply the cluster is free of supernova remnant material for of the time (Muno et al., 2006); however, we note that this estimated supernova rate is very uncertain. RSG winds inject cool material that can be ionised by an external radiation field, such as that found in a YMC (Mackey et al., 2014). This has already been demonstrated for W26, one of the four RSGs in Westerlund 1 (Wright et al., 2014; Mackey et al., 2015). Furthermore, the potential for the enriched RSG wind material to contribute to future star formation within the cluster depends on whether it is heated and ejected rapidly, or can remain confined, cool, and dense (Mackey et al., 2015).
@@
Recent JWST images of Westerlund 1 from the Extended Westerlund 1 and 2 Open Clusters Survey (EWOCS) show resolved shells and outflows from the four RSGs and other evolved cool massive stars in the cluster (Guarcello et al., 2025). The emission mechanism for gas and dust in Westerlund 1 is not clear, but its proximity to the core of such a massive cluster implies that the material was recently ejected from the cluster member stars. If the mid-IR emission is produced by dust, then the most likely source is winds and mass ejections from the cool supergiants and hypergiants in the cluster core.
@@
In this paper we explore, in a sense, the opposite case to that of Rogers & Pittard (2013), that of a cool RSG wind embedded in a hot, fast collective cluster wind from a YMC. We present a 3D HD simulation of our scenario to explore the efficiency of mass loading from evolved stellar winds, and use it to create synthetic dust-emission maps. There is some similarity between our work and previous simulations of the G2 cloud orbiting the Galactic Centre: De Colle et al. (2014) and Ballone et al. (2018) modelled a cool star emitting a slow wind and moving rapidly through the hot coronal gas around the supermassive black hole Sgr A∗. To our knowledge, no multi-dimensional simulations have been made of the interaction of stellar wind from a RSG embedded in a hot star-cluster wind, though such a situation must be common in YMCs.
@@
Our paper is organised as follows: In Sect. 2 we discuss our simulation setup and the values adopted for the RSG and cluster winds. In Sect. 3 we present the results of our simulation, including slices through the XY domain (Sect. 3.1), phase diagrams for the RSG wind and cluster wind material (Sect. 3.2), mass loading of the cluster wind over time (Sect. 3.3), synthetic dust continuum-emission maps for 11 m and 24 m derived from our simulation (Sect. 3.4), and how those maps correspond to observations from JWST (Sect. 3.5). In Sect. 4 we discuss the uncertainties in the winds of RSGs and clusters (Sect. 4.1) and in our modelling (Sect. 4.2), as well as the implications for this work in understanding the role of mass loading (Sect. 4.3), its connection to super star clusters (Sect. 4.4), and the need for follow-up observations (Sect. 4.5). We present our conclusions in Sect. 5, and a resolution study in Appendix A.
@@ 2.1 HD simulations
We performed our 3D Cartesian simulations using the (magneto-)HD code pion (Mackey et al., 2021). We employed the static mesh refinement option in pion to achieve high resolution only where required. For our radiative heating and cooling prescription, we used model 8 in pion as described in Green et al. (2019), which assumes a sufficiently intense extreme-ultraviolet (EUV) radiation field to keep H and He photoionised everywhere on the domain, and which includes three cooling sources. The first source is the maximum of the Solar metallicity collisional ionisation equilibrium curve described by Wiersma et al. (2009), and the cooling function for forbidden lines described by Henney et al. (2009). This ensures we account for the significant rates of cooling that occur in photoionisation equilibrium at K. The second source included is ionised hydrogen (Hummer, 1994) and helium (Rybicki & Lightman, 1979) Bremsstrahlung. The third source is the cooling rate of hydrogen recombination from Hummer (1994), where we assumed hydrogen is fully ionised. Heating is calculated by assuming each H recombination (excluding recombinations directly to the ground state, i.e. the Case B approximation) rapidly results in a re-ionisation by the intense Lyman-continuum radiation field found in YMCs, with associated photo-heating of 5 eV per photoionisation. We obtain an equilibrium gas temperature K for this prescription.
In Westerlund 1 the EUV radiation field is provided by the OB and WR stars in the cluster, which are observed to ionise the circumstellar nebulae of the cool stars to a high degree in the vicinity of the cluster (Andrews et al., 2018). This validates our above assumption that the medium is photoionised. Consistent with this observation, postprocessing of simulation snapshots in Sect. 3.4 with the torus radiation transfer code calculates the ionisation state of the gas as well as dust emission. From this we find that H is more than 99% ionised in the full domain except for within a radius of cm around the RSG where the wind is dense enough to self-shield from ionising radiation.
The simplified microphysics treatment is quite standard in HD modelling of nebulae (e.g. Garcia-Segura et al., 1996; Meyer et al., 2014; Mackey et al., 2025), where calculation of radiative transfer and non-equilibrium ionisation would make the simulations prohibitively expensive (Mathew et al., 2025). Haworth et al. (2015) investigated the differences between a simplified microphysics treatment and more complicated calculations with full radiative transfer and ionisation of H, He, and metals, applied to expanding H ii regions. They found that differences were predominantly driven by the equilibrium temperature of the photoionised gas, determined by the gas composition and spectrum of the radiation field.
@@
We performed our simulation using a 3D Cartesian domain of cm, using five static mesh refinement levels. At the coarsest level, , the entire domain is covered, and more refined levels with cover a sub-domain of diameter , i.e. each finer level has a domain two times smaller than the level preceding it. We included grid cells per level of refinement, such that the maximum grid resolution is cm.
@@
To simulate the stellar wind, we placed the RSG star at the origin, and the nested grids were centred on the star. Our aim is to study the ablation of the RSG wind, and thus a constant mass loss source term is applied, with a surface temperature K, a mass-loss rate = -1 (Mackey et al., 2015) and wind terminal velocity (Mauron & Josselin, 2011). The wind injection region is cm, or ten cells. This is much larger than the stellar radius, and thus we injected the wind at the terminal velocity.
@@
To simulate the cluster wind, we imposed a constant x-velocity km s-1, density g cm-3, and pressure dyn cm-2. While direct observational constraints are challenging, the cluster wind velocities cannot significantly exceed the adiabatic sound speed. This is of order 500-1000 km s-1, using the diffuse X-ray gas temperatures measured by Kavanagh et al. (2011). Our velocity is also consistent with that of Silich et al. (2011), but lower than theoretical predictions (e.g. Stevens & Hartwell, 2003; Härer et al., 2023), which are of order 2000-3000 km s-1. Our density value is based on the density of the X-ray emitting gas from Kavanagh et al. (2011). The upstream boundary condition was set to the fixed inflow conditions of the cluster wind, and the other external boundaries are set to zero-gradient (outflow).
@@ 3.1 XY slices
The ablation flow bears some similarity to the 2D von Karman vortex street, generated by subsonic flow past a solid object. Direct estimates of Reynolds numbers (Re) in simulations of this type are difficult due to the requirement for artificial viscosity. Since we are in between a laminar flow (Re ) and fully turbulent (Re s), we can assume our Re is of order s. With this assumption, for an idealised von Karman vortex street (where the characteristic length scale, is the diameter of the cylinder), the shedding frequency, is related to the Strouhal number (St) by the ratio of (in our case, twice the stand-off distance to the apex of the bow shock) and flow velocity . For flow past a cylinder is well defined and, importantly, subsonic, but in our case we have the complication of two different flow speeds (the cluster wind and the RSG wind) with widely differing sound speeds. The difficulty in assigning a characteristic velocity in such a multi-scale shear flow is addressed by Tan et al. (2021). If we nevertheless assume corresponds to the RSG wind velocity, we can calculate a characteristic timescale. As St only varies slightly for Re , we assumed St = 0.2 and find a typical shedding timescale of 3 kyr. However, due to the instability of the bow shock, is neither fixed nor constant over the bow shock, and so the vortices are only quasi-periodic. This is similar to the behaviour found by Wareing et al. (2007) in simulations of asymptotic giant branch stars, albeit with much lower relative velocities.
@@ 3.2 Phase diagrams
In Fig. 2 we show the density-temperature phase diagrams for only the RSG wind, only the cluster wind, and the sum of these components at the end of our simulation. We considered the RSG wind to be ‘cool’, material above the photoionisation equilibrium temperature K to be ‘mixed’, and material significantly hotter than this to be ‘hot’. In particular, RSG wind material that is over K must have been heated by interactions, shocks and other processes.
Focusing on the RSG wind first, there is a horizontal feature at low temperature that corresponds to material immediately surrounding the star, mainly unshocked stellar wind. At the bow shock, the RSG wind is ablated by the cluster wind and clumps break off. As seen in Fig. 1, these clumps can remain dense and cool. This explains the ‘hotspot’ around g cm-3 and the equilibrium temperature of K. As these clumps continue to be mixed, some of the material is heated and produces the long tail in that panel.
For the cluster wind in the centre panel of Fig. 2, there is a hotspot at the ambient cluster wind values of g cm-3 and K. As the cluster wind mixes with the RSG wind, some of this material is cooled. Much of the mixed material cools until it reaches the equilibrium temperature (K), and this is seen as a horizontal feature in the phase diagram. The presence of a feature at the equilibrium temperature in all three panels reflects the dynamical mixing of the two gas phases, which both heat the RSG wind and cool the cluster wind.
In Fig. 3 we show phase diagrams of versus temperature for the RSG wind material at 50 kyr intervals. There is a hotspot at the wind injection temperature, visible as the isolated point in the upper left in all panels. As the RSG wind material expands spherically, we see a vertical feature where the material reaches the equilibrium temperature with a dispersion in velocity at the bow shock, which widens over time as the simulation approaches a time-stationary state.
The RSG wind material is ablated by the cluster wind, and rapidly accelerated as it is heated and then exits the domain, and the effect increases with time. The similarity of the last two panels implies that the simulation is approaching a time-stationary state.
@@ 3.3 Mass loading
To get an idea of the efficiency of the mass loading over time, we compared the time evolution of ‘hot’ (K) and ‘cool’ (K) RSG wind material (see Fig. 4). As expected, the cool RSG wind mass increases rapidly at early times. It stays relatively constant over time for small radii, but appears to peak and then decrease for larger radii (Fig. 4a). As the contiguous tail is broken up and ablated, the cool material is removed from the domain more rapidly over time. The mass of hot RSG wind material changes rapidly at first (Fig. 4b), likely due to initial conditions. At later times it tends to increase to a quasi steady state for all radii. As this material is accelerated and exits the domain, we only find a lower limit for the amount of RSG material that is heated. This behaviour is consistent across different heating temperatures at small and large radii (Fig. 4c). The small fluctuations arise from the turbulent nature of the instabilities themselves.
The ratio of hot to cool RSG wind material tends to increase over time, again with turbulent fluctuations. Given that some of the heated wind is escaping from the domain, this represents a lower limit on the mass loading. Taking , this represents a mass-loading efficiency of 33% at late times, albeit a more detailed numerical study with a larger simulation domain is required to measure this number accurately. We study the numerical convergence of these results in Appendix A.
Second generation star formation from rapidly cooling shocked stellar winds has been studied in detail by Wünsch et al. (2017). However, that requires material to be shielded (i.e. not be photoionised), which is never the case in the central part of Westerlund 1 that we are considering due to the magnitude of the incident ionising radiation. We therefore do not expect there to be second generation star formation resulting from the ablation flows in this case.